OpenStudy (anonymous):

Help(:

OpenStudy (anonymous):

\[s _{26} for -31+(-27)+(-23)+(-19)+... enter the s u m\]

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

We are looking for, \(\large{S_{26}}\) And our initial term is \(a_{0}=-31\). So can we reasonably say that our problem is, \(\large{\sum\limits_{k=1}^{26}a_0+4,~a_0=-31}\) I believe this is the correct setup.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it like the last one where you add?

OpenStudy (wolfe8):

After you have the initial value(a0) and d(change), use the formula for sum.

OpenStudy (austinl):

Do you have a scientific calculator?

OpenStudy (jdoe0001):

http://speedcrunch.org/en_US/download.html <----

OpenStudy (wolfe8):

You can use this formula, which is something you should memorize.

OpenStudy (austinl):

\(a_0+4\) \(a_0=-31\) \(a_0+4=a_1=-27\) \(a_1+4=a_2=...\) And it continues on. I believe I made an error in the set up, but this is basically how you do it.

OpenStudy (wolfe8):

This page should help: http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html

OpenStudy (anonymous):

thank you(: